OpenStudy (cccmpt2):

How would I find one of the factors of 3x^2+x?

hartnn (hartnn):

factor out 'x' from 3x^2+x, what u get ?

OpenStudy (cccmpt2):

sorry, don't know?

OpenStudy (cccmpt2):

3^2+x?

hartnn (hartnn):

let me give you similar example, if it was 9x^3 + 4x, you would have gor x*(9x^2)+ x*(4) = x *(9x^2+4) got this ?

OpenStudy (anonymous):

x (3x+1) x=0 / 3x+1=0 x=0 / x=-1/3

hartnn (hartnn):

if you want to solve then, x=0 / x=-1/3 but if you just need factors, those are x and 3x+1